# In-Context Learning as Nonparametric Conditional Probability Estimation: Risk Bounds and Optimality

### v1

@@ Appendix D Proof of Lemma 12
Proof  We first establish a useful Lipschitz continuity property of the log-softmax function. This intermediate result is critical for bounding covering numbers in Lemma 12.
@@
Let $\mathbf{y}=(y_{1},\dots,y_{K})\in\mathbb{R}^{K}$, and define the softmax function $\mathbf{\Phi}:\mathbb{R}^{K}\to\Delta^{K-1}$ by
@@
where $\Delta^{K-1}$ is the probability simplex in $\mathbb{R}^{K}$. Then $\log\circ\mathbf{\Phi}$ is $K$-Lipschitz with respect to the $\|\cdot\|_{\infty}$-norm, i.e.,
@@
Proof  Consider the log-softmax function
$\log(\mathbf{\Phi}(\mathbf{y}))_{i}=\mathbf{y}_{i}-\log\left(\sum_{j=1}^{K}e^{\mathbf{y}_{j}}\right).$
Taking derivatives, for $i,k\in\{1,\dots,K\}$, we have
$\frac{\partial}{\partial\mathbf{y}_{k}}\log(\mathbf{\Phi}(\mathbf{y}))_{i}=\delta_{ik}-\mathbf{\Phi}(\mathbf{y})_{k},$
where $\delta_{ik}$ is the Kronecker delta. As $0\leq\mathbf{\Phi}(\mathbf{y})_{k}\leq 1$, the Jacobian’s entries $\delta_{ik}-\mathbf{\Phi}(\mathbf{y})_{k}$ are bounded by 1 in absolute value. Thus, the Lipschitz constant under the $\ell_{\infty}$-norm is at most $K$.
@@
$\log\mathcal{F}=\log(\mathbf{\Phi}(\mathcal{F}^{\prime})).$
The next lemma establish the covering number bound of $log\mathcal{F}$ with respect to the norm $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$.
@@
$\mathcal{N}\left(\delta,\log\mathcal{F},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\mathcal{N}\left(\tfrac{\delta}{2K},\mathcal{F}^{\prime},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right).$
Proof  Fix $\delta>0$. Let $\{\mathbf{f}_{j}\}_{j=1}^{J}$ be the centers of a minimal $\delta/(2K)$-cover of $\mathcal{F}^{\prime}$ under the empirical sup-norm $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$. That is, for every $\mathbf{f}\in\mathcal{F}^{\prime}$, there exists $j\in\{1,\dots,J\}$ such that
@@
Let $\mathcal{T}$ be a class of functions from $0,1]^{d}$ to $\mathbb{R}^{K}$, and let $\mathcal{F}_{\mathrm{id}}$ be a class of functions from $\mathbb{R}^{K}$ to $\mathbb{R}^{K}$, such that every $\mathbf{f}\in\mathcal{F}_{\mathrm{id}}$ is $L_{1}$-Lipschitz with respect to the sup-norm $\|\cdot\|_{\infty}$. Define the composed function class
@@
Proof  Fix $\delta>0$. Let $\widehat{\mathcal{T}}$ be a $\tfrac{\delta}{4KL_{1}}$-cover of $\mathcal{T}$ with respect to the empirical sup-norm $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$, and let $\widehat{\mathcal{F}}_{\mathrm{id}}$ be a $\tfrac{\delta}{4K}$-cover of $\mathcal{F}_{\mathrm{id}}$ with respect to the sup-norm $\|\cdot\|_{\infty}$. Define the composite approximation set
@@
Next we show that $\widehat{\mathcal{F}^{\prime}}$ is a $\tfrac{\delta}{2K}$-cover of $\mathcal{F}^{\prime}$ with respect to the empirical sup-norm $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$. Indeed, for any $\mathbf{f}\in\mathcal{F}_{\mathrm{id}}$ and $\boldsymbol{\phi}\in\mathcal{T}$, there exist $\hat{\mathbf{f}}\in\widehat{\mathcal{F}}_{\mathrm{id}}$ and $\hat{\boldsymbol{\phi}}\in\widehat{\mathcal{T}}$ such that
$\|\boldsymbol{\phi}-\hat{\boldsymbol{\phi}}\|_{\infty,\mathcal{Z}_{T}}\leq\tfrac{\delta}{4KL_{1}}\quad\text{and}\quad\|\mathbf{f}-\hat{\mathbf{f}}\|_{\infty}\leq\tfrac{\delta}{4K}.$
Then for each $t\in\{1,\ldots,T\}$, we have
@@
$\leq L_{1}\tfrac{\delta}{4KL_{1}}+\tfrac{\delta}{4K}=\tfrac{\delta}{2K}.$
@@
$\mathcal{N}\left(\delta,\log\mathcal{F},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)$ $\leq$ $\mathcal{N}\left(\tfrac{\delta}{2K},\mathcal{F}^{\prime},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)$
@@ Appendix E Proof of Lemma 13
Before presenting the detailed proof of Lemma 13, we first introduce the notion of pseudo-dimension which will be used in the proof of Lemma  13.
@@
Proof [Proof of Lemma 13.] Our proof for Lemma 13 is similar to that of Lemma 30 in Jiao et al. (2024), Lemma 8 in Gurevych et al. (2022), and Theorem 6 in Bartlett et al. (2019). The main idea is to bound the pseudo-dimension of subsets of $\mathcal{T}$ and then utilize well-known results developed by Anthony and Bartlett (2009) to bound the covering entropy by pseudo-dimensions. Here $\mathcal{T}=\mathcal{T}(M,H,D,W_{1},B_{1},S_{1},\gamma)$ defined in (3.3) in the main text.
@@
$\mathcal{N}(\delta_{1},\mathcal{G},\|\cdot\|_{\infty,\mathcal{Z}_{T}})\leq\mathcal{N}(\delta_{1},\mathcal{B}_{B_{1}},\|\cdot\|_{\infty})\leq\left(\frac{3B_{1}}{\delta_{1}}\right)^{T}\leq\left(\frac{3B_{1}}{\delta_{1}}\right)^{{\rm Pdim}(\mathcal{G})},\ \text{for}\ T<{\rm Pdim}(\mathcal{G}).$
@@
In the end of this subsection, we present some lemmas that are used in the proof of Lemma  13.
@@
Let $\mathcal{F}$ be a class of real-valued functions defined on a domain $\mathcal{X}$ with range bounded in the interval $[0,B]$. Let $\epsilon>0$, and suppose that the pseudo-dimension of $\mathcal{F}$ is $d$. Then the $\ell_{\infty}$ covering number satisfies
@@
which is further bounded by
$\left(\frac{emB}{\epsilon d}\right)^{d}\quad\text{for }m\geq d.$
@@ Appendix F Proof of Lemma 14
$\leq$ $\sum_{k=1}^{L}\|\mathcal{B}_{k+1}(\mathbf{g})\|_{Lip}\|(W^{(k)}\cdot+b^{(k)})\circ\mathcal{A}_{k}(\mathbf{f})(\mathbf{x})-({W^{{}^{\prime}(k)}}\cdot+{b^{{}^{\prime}(k)}})\circ\mathcal{A}_{k}(\mathbf{f})(\mathbf{x})\|_{\infty}.$
Next, we explicitly derive the following uniform bound for each intermediate mapping. According the definition of $\mathcal{F}_{\rm id}(L,W,S)$ given in (5), we have
@@
Similar to the arguments for Lemma  13, we fix the positions of these nonzero parameters of the network $F_{id}(L,W,S)$ and then its covering number with respect to the $\|\cdot\|_{\infty}$-norm is bounded by
@@ Appendix G Proof of Lemma 16
The Kullback–Leibler divergence and the Hellinger distance are two classical measures of discrepancy between probability distributions; see Tsybakov (2008, Section 2.4) and Birgé and Massart (1998, Lemma 5). These two measures are closely related and satisfy several useful inequalities, which are fundamental in nonparametric estimation theory. Recall that the squared Hellinger distance between $P$ and $Q$ is defined as
$H^{2}(P,Q)=\left(1-\int\sqrt{dP\,dQ}\right),$
as given in Birgé and Massart (1998, Section 7.3).
We proceed to establish the inequality stated in Lemma 16.
Proof [Proof of Lemma 16] We first show that $H^{2}(P,Q)\leq\frac{1}{2}\mathrm{KL}_{2}(P,Q)\leq\frac{1}{2}\mathrm{KL}_{B}(P,Q)$. As the second inequality follows from the definition of the truncated KL-divergence, we only show $H^{2}(P,Q)\leq\frac{1}{2}\mathrm{KL}_{2}(P,Q)$. Let $P=P^{a}+P^{s}$ be the Lebesgue decomposition of $P$ with respect to $Q$, where $P^{a}\ll Q$ and $P^{s}\perp Q$. Let $\mu=\frac{1}{2}(P+Q)$ be the dominating measure and let $p,p^{a},p^{s},q$ denote the corresponding densities of $P,P^{a},P^{s},Q$ with respect to $\mu$, respectively. Since $p^{s}q=0$, it follows that
@@
$\mathrm{KL}_{2}(P,Q)$ $=$ $\int\left(2\wedge\log\frac{dP}{dQ}\right)dP$
@@
Thus, we obtain $H^{2}(P,Q)\leq\frac{1}{2}KL_{2}(P,Q)$.
The proof of the upper bound on $\mathrm{KL}_{B}(P,Q)$ in terms of $H^{2}(P,Q)$ follows a similar approach to that of Theorem 5 of Wong and Shen (1995) and Lemma 3.4 of Bos and Schmidt-Hieber (2022). For completeness, we also provide a detailed derivation. Note that by second order Taylor expansion, we have $x\log(x)\leq x-1+\frac{1}{2}(x-1)^{2}/(x\wedge 1)$ for $x>0$. Consequently,
@@
$\mathrm{KL}_{B}(P,Q)$ $=$ $\int_{p^{a}/q\leq e^{B}}\frac{p^{a}}{q}\log\left(\frac{p^{a}}{q}\right)qd\mu+B\int_{\frac{dP}{dQ}>e^{B}}dP$
@@
$\mathrm{KL}_{B}(P,Q)\leq 2\left(e^{B/2}+B(1-e^{-B/2})^{-2}\right)H^{2}(P,Q).$
@@
$\mathrm{KL}_{B}(P,Q)\leq$ $\ 2\int_{p^{a}/q\leq e^{B}}(p^{a}-\sqrt{p^{a}q})d\mu+2e^{B/2}H^{2}(P,Q)+B\int_{\frac{dP}{dQ}>e^{B}}dP$
@@
Thus, we obtain $\mathrm{KL}_{B}(P,Q)\leq 10e^{B/2}H^{2}(P,Q)$. Consequently,
$\frac{1}{2}\mathrm{KL}_{B}(P,Q)\leq 5e^{B/2}H^{2}(P,Q).$
@@ Appendix H Proof of Lemma 19
Proof [Proof of Lemma 19] Let $\mathcal{F}_{\mathrm{MLP}}(L,W,S)$ be defined as in (9), where $\boldsymbol{\Phi}$ denotes the softmax mapping. For any empirical input set $\mathcal{Z}_{T}\subset[0,1]^{d}$, we aim to bound the covering number of the log-transformed class $\log\big{(}\mathcal{F}_{\mathrm{MLP}}(L,W,S)\big{)}$ under the empirical sup-norm $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$.
By Lemma 31, for any $\delta>0$, we have
$\mathcal{N}\left(\delta,\log\left(\mathcal{F}_{\mathrm{MLP}}(L,W,S)\right),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\mathcal{N}\left(\tfrac{\delta}{2K},\mathcal{F}_{\mathrm{id}}(L,W,S),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right).$
Setting $\delta=\tfrac{1}{T}$, it follows that
$\mathcal{N}\left(\tfrac{1}{T},\log\left(\mathcal{F}_{\mathrm{MLP}}(L,W,S)\right),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\mathcal{N}\left(\tfrac{1}{2KT},\mathcal{F}_{\mathrm{id}}(L,W,S),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right).$
Applying the standard pseudo-dimension-based covering bound (see Lemma 36, Theorem 12.2 of Anthony and Bartlett (2009)), we obtain
$\mathcal{N}\left(\tfrac{1}{2KT},\mathcal{F}_{\mathrm{id}}(L,W,S),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\left(\frac{2eKC\log T\cdot T^{2}}{\operatorname{Pdim}(\mathcal{F}_{\mathrm{id}})}\right)^{\operatorname{Pdim}(\mathcal{F}_{\mathrm{id}})},$
where the output range of $\mathcal{F}_{\mathrm{id}}$ is bounded by $C\log T$ for some universal constant $C>0$.
@@
$\underline{c}\cdot SL\log(S/L)\leq\operatorname{Pdim}(\mathcal{F}_{\mathrm{id}}(L,W,S))\leq\overline{c}\cdot SL\log S.$
@@
$\mathcal{V}_{\infty}\left(\tfrac{1}{T},\log\left(\mathcal{F}_{\mathrm{MLP}}(L,W,S)\right),T\right)\leq C\cdot SL\log S\cdot\log T,$

### v2

@@ Appendix D Proof of Lemma 12
Proof [Proof of Lemma 12] We first establish a useful Lipschitz continuity property of the log-softmax function. This intermediate result is critical for bounding covering numbers in Lemma 12.
@@
Let $\mathbf{y}=(y_{1},\dots,y_{K})^{\top}\in\mathbb{R}^{K}$, and define the softmax function $\mathbf{\Phi}:\mathbb{R}^{K}\to\mathcal{S}^{K}$ by
@@
where $\mathcal{S}^{K}$ is the probability simplex in $\mathbb{R}^{K}$. Then $\log\circ\mathbf{\Phi}$ is $K$-Lipschitz with respect to the $\|\cdot\|_{\infty}$-norm, i.e.,
@@
Proof  Consider the log-softmax function $\log(\mathbf{\Phi}(\mathbf{y}))_{i}=y_{i}-\log\left(\sum_{j=1}^{K}e^{y_{j}}\right)$. Taking derivatives, we have for $i,k\in\{1,\dots,K\}$,
$\frac{\partial}{\partial y_{k}}\log(\mathbf{\Phi}(\mathbf{y}))_{i}=\delta_{ik}-\mathbf{\Phi}(\mathbf{y})_{k},$
where $\delta_{ik}$ is the Kronecker delta. As $0\leq\mathbf{\Phi}(\mathbf{y})_{k}\leq 1$, the Jacobian’s entries $\delta_{ik}-\mathbf{\Phi}(\mathbf{y})_{k}$ are bounded by 1 in absolute value. Thus, the Lipschitz constant is at most $K$.
@@
$\log(\mathcal{F})=\log(\mathbf{\Phi}(\mathcal{F}^{\prime})).$
The next lemma establish the covering number bound of $\log\mathcal{F}$ with respect to the norm $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$.
@@
$\mathcal{N}\left(\delta,\log(\mathcal{F}),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\mathcal{N}\left(\tfrac{\delta}{2K},\mathcal{F}^{\prime},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right).$
Proof  Fix $\delta>0$. Let $\{\mathbf{f}_{j}\}_{j=1}^{J}$ be the centers of a minimal $\delta/(2K)$-cover of $\mathcal{F}^{\prime}$ with respect to $\|\cdot\|_{\mathcal{Z}_{T}}$. That is, for every $\mathbf{f}\in\mathcal{F}^{\prime}$, there exists $j\in\{1,\dots,J\}$ such that
@@
Let $\mathcal{T}$ be a class of functions from $0,1]^{d}$ to $\mathbb{R}^{K}$, and let $\mathcal{F}_{\mathrm{id}}$ be a class of functions from $\mathbb{R}^{K}$ to $\mathbb{R}^{K}$, such that every $\mathbf{f}\in\mathcal{F}_{\mathrm{id}}$ is $L_{1}$-Lipschitz with respect to $\|\cdot\|_{\infty}$. Define the composed function class
@@
Proof  Fix $\delta>0$. Let $\widehat{\mathcal{T}}$ be a $\tfrac{\delta}{4KL_{1}}$-cover of $\mathcal{T}$ with respect to $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$, and let $\widehat{\mathcal{F}}_{\mathrm{id}}$ be a $\tfrac{\delta}{4K}$-cover of $\mathcal{F}_{\mathrm{id}}$ with respect to $\|\cdot\|_{\infty}$. Define the composite approximation set
@@
Next, we show that $\widehat{\mathcal{F}^{\prime}}$ is a $\tfrac{\delta}{2K}$-cover of $\mathcal{F}^{\prime}$ with respect to $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$. Indeed, for any $\mathbf{f}\in\mathcal{F}_{\mathrm{id}}$ and $\boldsymbol{\phi}\in\mathcal{T}$, there exist $\hat{\mathbf{f}}\in\widehat{\mathcal{F}}_{\mathrm{id}}$ and $\hat{\boldsymbol{\phi}}\in\widehat{\mathcal{T}}$ such that Then for each $t\in\{1,\ldots,T\}$, we have
@@
$\leq L_{1}\tfrac{\delta}{4KL_{1}}+\tfrac{\delta}{4K}$
$=\tfrac{\delta}{2K}.$
@@
$\mathcal{N}\left(\delta,\log(\mathcal{F}),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)$ $\leq$ $\mathcal{N}\left(\tfrac{\delta}{2K},\mathcal{F}^{\prime},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)$
@@ Appendix E Proof of Lemma 13
Before presenting the detailed proof of Lemma 13, we first introduce the notion of pseudo-dimension which will be used in the proof of Lemma 13.
@@
Proof [Proof of Lemma 13.] Our proof for Lemma 13 is similar to that of Lemma 30 in Jiao et al. (2024), Lemma 8 in Gurevych et al. (2022), and Theorem 6 in Bartlett et al. (2019). The main idea is to bound the pseudo-dimension of subsets of $\mathcal{T}$ and then utilize well-known results developed by Anthony and Bartlett (2009) to bound the entropy by pseudo-dimensions. Here $\mathcal{T}=\mathcal{T}(M,H,D,W_{1},B_{1},S_{1},\gamma)$ defined in (3.3) in the main text.
@@
$\mathcal{N}(\delta_{1},\mathcal{G},\|\cdot\|_{\infty,\mathcal{Z}_{T}})\leq\mathcal{N}(\delta_{1},\mathcal{B}_{B_{1}},\|\cdot\|_{\infty})\leq\left(\tfrac{3B_{1}}{\delta_{1}}\right)^{T}\leq\left(\tfrac{3B_{1}}{\delta_{1}}\right)^{\mathrm{Pdim}(\mathcal{G})},\qquad\,T<\mathrm{Pdim}(\mathcal{G})\,.$
@@
In the end of this subsection, we present some lemmas that are used in the proof of Lemma 13.
@@
Let $\mathcal{F}$ denote a class of real-valued functions on a domain $\mathcal{X}$, with range in $[0,B]$. Let $\epsilon>0$, and assume $\mathcal{F}$ has pseudo-dimension $d$. Then
@@
which is further bounded by $\left(\frac{emB}{\epsilon d}\right)^{d}$ for $m\geq d$.
@@ Appendix F Proof of Lemma 14
$\leq$ $\sum_{k=1}^{L}\|\mathcal{B}_{k+1}(\mathbf{g})\|_{Lip}\cdot\|(W^{(k)}\cdot+b^{(k)})\circ\mathcal{A}_{k}(\mathbf{f})(\mathbf{x})-({W^{{}^{\prime}(k)}}\cdot+{b^{{}^{\prime}(k)}})\circ\mathcal{A}_{k}(\mathbf{f})(\mathbf{x})\|_{\infty},$
where $\|\mathcal{B}_{k+1}(\mathbf{g})\|_{Lip}$ denotes the corresponding Lipschitz constant of $\mathcal{B}_{k+1}(\mathbf{g})$. Next, we explicitly derive the following uniform bound for each of the intermediate mappings. According to the definition of $\mathcal{F}_{\rm id}(L,W,S)$ provided in (5), we obtain
@@
Similar to the arguments for Lemma 13 in Appendix E, we fix the positions of these nonzero parameters of the network $F_{id}(L,W,S)$ and then its covering number with respect to the $\|\cdot\|_{\infty}$-norm is bounded by
@@ Appendix G Proof of Lemma 16
The Kullback-Leibler divergence and Hellinger distance are two classical measures of discrepancy between probability distributions; see Tsybakov (2008, Section 2.4) and Birgé and Massart (1998, Lemma 5). These two measures are closely related and satisfy several useful inequalities, which are fundamental in nonparametric estimation theory. Recall that the squared Hellinger distance between $P$ and $Q$ is defined as $H^{2}(P,Q)=1-\int\sqrt{dP\,dQ}$ (Birgé and Massart, 1998, Section 7.3). We proceed to establish the inequality stated in Lemma 16.
Proof [Proof of Lemma 16] We first show that $H^{2}(P,Q)\leq\frac{1}{2}\mathrm{KL}_{2}(P\|Q)\leq\frac{1}{2}\mathrm{KL}_{B}(P\|Q)$. As the second inequality follows from the definition of the truncated KL-divergence, we only show $H^{2}(P,Q)\leq\frac{1}{2}\mathrm{KL}_{2}(P\|Q)$. Let $P=P^{a}+P^{s}$ be the Lebesgue decomposition of $P$ with respect to $Q$, where $P^{a}\ll Q$ and $P^{s}\perp Q$. Let $\mu=\frac{1}{2}(P+Q)$ be the dominating measure and let $p,p^{a},p^{s},q$ denote the corresponding densities of $P,P^{a},P^{s},Q$ with respect to $\mu$, respectively. Since $p^{s}q=0$, it follows that
@@
$\mathrm{KL}_{2}(P\|Q)$ $=$ $\int\left(2\wedge\log\frac{dP}{dQ}\right)dP$
@@
Therefore, we obtain $H^{2}(P,Q)\leq\frac{1}{2}KL_{2}(P\|Q)$.
The proof of the upper bound on $\mathrm{KL}_{B}(P\|Q)$ in terms of $H^{2}(P,Q)$ follows a similar approach to that of Theorem 5 of Wong and Shen (1995) and Lemma 3.4 of Bos and Schmidt-Hieber (2022). For completeness, we also provide a detailed derivation. Note that by second order Taylor expansion, we have $x\log(x)\leq x-1+\frac{1}{2}(x-1)^{2}/(x\wedge 1)$ for $x>0$. Consequently,
@@
$\mathrm{KL}_{B}(P\|Q)$ $=$ $\int_{p^{a}/q\leq e^{B}}\frac{p^{a}}{q}\log\left(\frac{p^{a}}{q}\right)qd\mu+B\int_{\frac{dP}{dQ}>e^{B}}dP$
@@
$\mathrm{KL}_{B}(P\|Q)\leq 2\left(e^{B/2}+B(1-e^{-B/2})^{-2}\right)H^{2}(P,Q).$
@@
$\mathrm{KL}_{B}(P\|Q)\leq$ $\ 2\int_{p^{a}/q\leq e^{B}}(p^{a}-\sqrt{p^{a}q})d\mu+2e^{B/2}H^{2}(P,Q)+B\int_{\frac{dP}{dQ}>e^{B}}dP$
@@
Therefore, we obtain $\mathrm{KL}_{B}(P\|Q)\leq 10e^{B/2}H^{2}(P,Q)$. Consequently,
$\frac{1}{2}\mathrm{KL}_{B}(P\|Q)\leq 5e^{B/2}H^{2}(P,Q).$
@@ Appendix H Proof of Lemma 21
Proof [Proof of Lemma 21] Recall that $\mathcal{F}_{\rm mlp}(L,W,S)=\mathbf{\Phi}(\mathcal{F}_{\mathrm{id}}(L,W,S))$, it follows from Lemma 34 that
$\mathcal{N}\left(\frac{1}{T},\log\left(\mathcal{F}_{\mathrm{mlp}}(L,W,S)\right),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\mathcal{N}\left(\frac{1}{2KT},\mathcal{F}_{\mathrm{id}}(L,W,S),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right).$
Consider the function classes
${\mathcal{F}^{k}_{\mathrm{id}}(L,W,S)}^{+}$ $=$ $\{\sigma(\mathbf{e}_{k}^{\top}\mathbf{G}):\mathbf{G}=(G_{1},\dots,G_{K})^{\top}\in\mathcal{F}_{\mathrm{id}}(L,W,S)\},$
${\mathcal{F}^{k}_{\mathrm{id}}(L,W,S)}^{-}$ $=$ $\{\sigma(-\mathbf{e}_{k}^{\top}\mathbf{G}):\mathbf{G}=(G_{1},\dots,G_{K})^{\top}\in\mathcal{F}_{\mathrm{id}}(L,W,S)\},$
where $\sigma(x)=\max(x,0)$ is the ReLU activation function and $\mathbf{e}_{k}=(0,\dots,0,1,0,\dots,0)^{\top}\in\mathbb{R}^{K}$ with $1$ lying in the $k$-th component of $\mathbf{e}_{k}$. It is readily seen that both $\mathcal{F}^{k}_{\mathrm{id}}(L,W,S)$$^{+}$ and $\mathcal{F}^{k}_{\mathrm{id}}(L,W,S)$$^{-}$ consist of ReLU networks with depth $L+1$, width $W$, and size $S+1$. Since $\|\mathbf{G}\|_{\infty}\leq B_{2}$ for any $\mathbf{G}\in\mathcal{F}_{\mathrm{id}}(L,W,S)$, it follows that $|\sigma(\mathbf{e}_{k}^{\top}\mathbf{G})|\leq B_{2}$ and $|\sigma(-\mathbf{e}_{k}^{\top}\mathbf{G})|\leq B_{2}$. Therefore, by Lemma 39 (Theorem 12.2 of Anthony and Bartlett (2009)), we obtain
$\mathcal{N}\left(\frac{1}{4KT},{\mathcal{F}^{k}_{\mathrm{id}}(L,W,S)}^{+},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)$ $\leq$ $\left(\frac{4eKB_{2}T^{2}}{\mathrm{Pdim}({\mathcal{F}^{k}_{\mathrm{id}}(L,W,S)}^{+})}\right)^{\mathrm{Pdim}({\mathcal{F}^{k}_{\mathrm{id}}(L,W,S)}^{+})},$
@@
$\underline{c}(S+1)(L+1)\log\frac{S+1}{L+1}\leq\mathrm{Pdim}({\mathcal{F}^{k}_{\mathrm{id}}(L,W,S)}^{+})\leq\overline{c}(S+1)(L+1)\log(S+1).$
Consequently,
$\mathcal{N}\left(\frac{1}{4KT},{\mathcal{F}^{k}_{\mathrm{id}}(L,W,S)}^{+},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\left(\frac{4eKB_{2}T^{2}}{\underline{c}(S+1)(L+1)\log\frac{S+1}{L+1}}\right)^{\overline{c}(S+1)(L+1)\log(S+1)},$
Similarly, we have
$\mathcal{N}\left(\frac{1}{4KT},{\mathcal{F}^{k}_{\mathrm{id}}(L,W,S)}^{-},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\left(\frac{4eKB_{2}T^{2}}{\underline{c}(S+1)(L+1)\log\frac{S+1}{L+1}}\right)^{\overline{c}(S+1)(L+1)\log(S+1)},$
Also note that $G_{k}=\sigma(\mathbf{e}_{k}^{\top}\mathbf{G})-\sigma(-\mathbf{e}_{k}^{\top}\mathbf{G})$, it follows that
$\mathcal{N}\left(\frac{1}{2KT},{\mathcal{F}_{\mathrm{id}}(L,W,S)},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\right)\leq\left(\frac{4eKB_{2}T^{2}}{\underline{c}(S+1)(L+1)\log\frac{S+1}{L+1}}\right)^{2\overline{c}K(S+1)(L+1)\log(S+1)}.$
@@
$\mathcal{V}_{\infty}\left(\frac{1}{T},\log\left(\mathcal{F}_{\mathrm{mlp}}(L,W,S)\right),T\right)\leq CKSL\log(S)\log(B_{2}KT),$